# Infinitary combinatorics in condensed mathematics and strong homology

### v1

@@ 4. Products of compact projectives
Recall the introductory account of condensed mathematics concluding Section 2.1 above. For motivation, as is standard, we began by emphasizing the shortcomings of, for example, the category of Hausdorff topological abelian groups, together with its faithful embedding into the abelian category $\mathsf{Cond(Ab)}$ as one far-reaching kind of remedy. Note, though, that the provision of an abelian — and even a symmetric monoidal abelian — category structure extending that of some ill-behaved preabelian category isn’t, in and of itself, particularly novel: Yoneda embeddings into abelian sheaf categories will always have these features (cf. comments, e.g., at [73, p. 9]). The distinction of condensed categories begins with something further, namely the conjunction of such features with the existence of a class of compact projective generators; we record the relevant definitions just below. Let us more immediately recall, though, that this class derives from the existence in the underlying category ProFin of the condensed site of a generating class of projective objects, and that the latter, as noted in Section 2.1 above, are precisely the extremally disconnected compact Hausdorff spaces, or $\mathsf{ED}$ spaces, for short. Note for future reference that the $\mathsf{ED}$ spaces also admit characterization as the retracts of the Čech–Stone compactifications $\beta X$ of discrete spaces $X$ (these $\beta X$ are accordingly termed the free objects of $\mathsf{CHaus}$ in [69]); recall as well the description of any such $\beta X$ as ($(\{\mathcal{U}\subseteq P(X)\mid\mathcal{U}\textnormal{ is an ultrafilter on }X\},\tau)$, where the topology $\tau$ on $\beta X$ is generated by the sets $N_{I}:=\{\mathcal{U}\in\beta X\mid I\in\mathcal{U}\}$ for $I\subseteq X$.
@@
    If $\mathcal{C}=\mathsf{Cond(Ab)}$ is the category of condensed abelian groups, then $\mathcal{C}^{\mathrm{cp}}$ is the category of direct summands of $\mathbb{Z}[S]$ for extremally disconnected $S$, which generates $\mathcal{C}$ under small colimits. [73, pp. 74-75]
@@
This proposition traces at the latest to Walter Rudin around 1960, by way of the attribution in [30], and induces, in turn, a number of effects and questions within the field of condensed mathematics. By item (4) above, for example, it implies that $\mathsf{Cond(Set)}^{\mathrm{cp}}$ is not closed under products. Subtler questions attach to item (5). There, $\mathbb{Z}[-]:\mathsf{Cond(Set)}\to\mathsf{Cond(Ab)}$ denotes the left adjoint to the forgetful functor $\mathsf{Cond(Ab)}\to\mathsf{Cond(Set)}$, and within $\mathsf{Cond(Ab)}$, the relation $\mathbb{Z}[S]\otimes\mathbb{Z}[T]=\mathbb{Z}[S\times T]$ renders $\mathbb{Z}[-]$-images of products of $\mathsf{ED}$ spaces computationally ubiquitous. These are not, in general, projective, although the main result in this direction requires more argument than one might at first expect; the proof given in [23], for example, answers an open question from [4] in Banach space theory [^10] That question, of whether the Banach space $C(\beta\mathbb{N}\times\beta\mathbb{N})$ is separably injective, typifies how this stratum of projectivity questions dualizes to injectivity questions in Banach settings; see [23, Prop. 3.17]. along the way:
@@
See [23, Appendix to III and pp. 26–27] for further discussion of these issues; the latter includes, for example, the following observation: by item (5), $\mathsf{Cond(Ab)}$ is rich in projective objects, i.e., in $A=\mathbb{Z}[S]$ for which $\mathrm{Ext}^{i}_{\mathsf{Cond(Ab)}}(A,-)=0$ for all $i>0$. For such to be internally projective, though, would entail that the internal Ext functors
@@
For any field $K$ and $\mathsf{ED}$ space $S$, the constant sheaf $\mathcal{K}$ on $S$ is of injective dimension at most one; it is zero if $K$ is finite. In contrast, if $\mathsf{ED}$ spaces $S$ and $T$ are each Čech-Stone compactifications of sets of cardinality at least $\aleph_{\omega}$, then the injective dimension of the constant sheaf $\mathcal{K}$ on their product is infinite.
@@ 4.1. Sheaves on $\mathsf{ED}$ spaces and their products
We recall a few further fundamental points before turning to our main argument; for a fuller review, see, for example, [45, Chapter 2].
@@
Recall that a sheaf $\mathcal{G}$ on $X$ is fully determined by its stalks $\mathcal{G}_{x}$ $(x\in X)$, so that the sheafification of a presheaf $\mathcal{F}$ may be defined simply as that sheaf $\widetilde{\mathcal{F}}$ on $X$ with ~$\widetilde{\mathcal{F}}_{x}=\mathcal{F}_{x}$ for all $x\in X$.
@@
with the left-hand side computed, of course, in the category concluding the previous sentence [45, 2.7.5]. Therein, injective objects admit the standard definition; there are enough of them, and the injective dimension $\mathrm{injdim}_{X}(\mathcal{F})$ of any sheaf $\mathcal{F}$ is the minimal length of an injective resolution of $\mathcal{F}$.
@@
of $\pi_{*}\mathcal{K}$ over $s$. By the compactness of $T$, for every locally constant $f:U\times T\to K$ there exists a neighborhood $V$ of $s$ and locally constant $g:T\to K$ such that $f(s^{\prime},t)=g(t)$ for all $(s^{\prime},t)\in V\times T$. It follows that $(\pi_{*}\mathcal{K})_{s}$ equals the $K$-vector space of continuous functions $T\to K$ (with discrete codomain); moreover, the latter is isomorphic to $W=\bigoplus_{w(T)}K$ by the argument of [36, Corollary 97.7]. The lemma’s concluding claim is now immediate: since the natural map from the constant sheaf $\mathcal{W}$ on $S$ to $\pi_{*}\mathcal{K}$ induces isomorphisms of stalks, it is an isomorphism of sheaves.
@@
for any $q\geq 0$, where $\iota^{*}$ denotes the pullback of the inclusion $\pi^{-1}(s)\hookrightarrow S\times T$ [63, Theorem 17.2]. Since pullback functors are exact and $\mathrm{H}^{q}(T;-)=0$ for any $q>0$ and profinite $T$ [84, Theorem 5.1], this in turn implies that $\pi_{*}$ is exact. The lemma’s first claim now follows from the general fact that pushforwards preserve injectives [45, II.4.1.3], since this, when combined with the foregoing, implies that the image $\pi_{*}\mathcal{J}$ of a finite injective resolution $\mathcal{J}$ of $\mathcal{K}$ is a finite injective resolution of $\pi_{*}\mathcal{K}$. ∎
@@
For every field $K$ and topological space $S$ and $K$-vector space $W$, the following conditions are equivalent:
@@
$(1)\Rightarrow(2)$: Much as in the argument of the preceding lemma, the point is that the pullback functor $\iota^{*}$ associated to the inclusion $\iota:U\hookrightarrow S$ is both exact and preserves injectives [45, II.6.10]. Hence $\iota^{*}$ applied to a length-$n$ injective resolution of the constant sheaf $\mathcal{W}$ over $S$ is a length-$n$ injective resolution of the constant sheaf $\mathcal{W}=\iota^{*}\mathcal{W}$ over $U$; this implies (2).
$(2)\Rightarrow(1)$: First we establish the following:
Claim 4.9.
A sheaf $\mathcal{F}$ of $K$-vector spaces is injective if and only if $\mathrm{Ext}_{K}^{1}(\iota_{!}\mathcal{K},\mathcal{F})=0$ for every inclusion $\iota$ of an open subset $U$ into $S$.
Proof.
The forward implication is standard. For the reverse implication, suppose that $\mathrm{Ext}_{K}^{1}$ vanishes in the manner described and let $f:\mathcal{A}\to\mathcal{F}$ and $g:\mathcal{A}\hookrightarrow\mathcal{B}$ be morphisms of sheaves over $S$; we need to show that $f$ extends to an $h:\mathcal{B}\to\mathcal{F}$. To that end, write $\mathcal{A}^{\prime}\leq\mathcal{A}^{\prime\prime}$ if $\mathcal{A}^{\prime}$ is a subsheaf of $\mathcal{A}^{\prime\prime}$, regard $\mathcal{A}$ as a subsheaf of $\mathcal{B}$, and consider the natural partial ordering of pairs $(\mathcal{A}^{\prime},h^{\prime})$ for which $\mathcal{A}\leq\mathcal{A}^{\prime}\leq\mathcal{B}$ and $h^{\prime}:\mathcal{A}^{\prime}\to\mathcal{F}$ satisfies $f=h^{\prime}\circ g$. Since this order satisfies the conditions of Zorn’s Lemma, our task reduces to showing that if there exists for some such $(\mathcal{A}^{\prime},h^{\prime})$ an $x\in S$ with $\mathcal{A}_{x}^{\prime}\lneq\mathcal{B}_{x}$ then there exists an $\mathcal{A}^{\prime\prime}$ with $\mathcal{A}^{\prime}\lneq\mathcal{A}^{\prime\prime}\leq\mathcal{B}$ and an $h^{\prime\prime}:\mathcal{A}^{\prime\prime}\to\mathcal{F}$ properly extending $h^{\prime}$.
To that end, fix such an $x$ and a nonzero $b\in\mathcal{B}_{x}\backslash\mathcal{A}_{x}^{\prime}$ and a representative $s\in\mathcal{B}(U)$ of $b$ for some open $U\in S$ containing $x$. Observe (by checking stalks) that $s$ generates a sheaf on $U$ which is isomorphic to, and which we are therefore justified in identifying with, $\mathcal{K}$. Let $\iota$ denote the inclusion $U\to S$ and let $\mathcal{A}^{\prime\prime}=\mathcal{A}^{\prime}+\iota_{!}\mathcal{K}$ and apply $\mathrm{Hom}_{K}(-,\mathcal{F})$ to the short exact sequence
$0\to\mathcal{A}^{\prime}\to\mathcal{A}^{\prime}+\iota_{!}\mathcal{K}\to\iota_{!}\mathcal{K}\to 0$
of sheaves on $S$. Our premise, applied to the fragment of the induced long exact sequence
[…]
implies that $e$ is surjective, and hence that $h^{\prime}$ extends to an an element $h^{\prime\prime}:\mathcal{A}^{\prime}+\iota_{!}\mathcal{K}\to\mathcal{F}$ of $e^{-1}(h^{\prime})$. This $h^{\prime\prime}$ and $\mathcal{A}^{\prime\prime}=\mathcal{A}^{\prime}+\iota_{!}\mathcal{K}$ conclude the argument. ∎
Fix now an $n$ as in (2) and a resolution
$0\to\mathcal{W}\to\mathcal{J}^{0}\to\cdots\to\mathcal{J}^{n}\to\mathcal{F}\to 0$
@@
of the sheaf $\mathcal{W}$ over $S$ in which each of the $\mathcal{J}^{i}$ are injective. By assumption, for every inclusion $\iota$ of an open subset $U$ into $S$,
$0=\mathrm{H}^{n+1}(U;\mathcal{W})$ $=\mathrm{H}^{n+1}(U;\iota^{*}\mathcal{W})$
$=\mathrm{Ext}_{K}^{n+1}(\mathcal{K},\iota^{*}\mathcal{W})$
$=\mathrm{Ext}_{K}^{n+1}(\iota_{!}\mathcal{K},\mathcal{W})$
$=\mathrm{Ext}_{K}^{1}(\iota_{!}\mathcal{K},\mathcal{F}),$
@@
hence by Claim 4.9, $\mathcal{F}$ is injective, completing the argument. Here the penultimate equality follows from the $\iota_{!}$$\iota^{*}$ adjunction together with the fact that such $\iota^{*}$ are exact and preserve injectives [45, II.6.9(ii)], and the last equality is via a standard dimension-shifting argument, such as appears, for example, in [83, Exercise 2.4.3]. ∎
@@
Against this background, we turn to the groups $\mathrm{H}^{q}(U;\mathcal{W})$ for such $W$ and open subsets $U\subseteq S$ as appear in Lemmas 4.7 and 4.8 above; since such $U$ admit a convenient characterization when $S$ is the Čech-Stone compactification $\beta X$ of some discrete space $X$, we will focus our attention on $S$ of this sort. Note that there is no loss of generality in doing so, since every $\mathsf{ED}$ space $T$ is a retract of some such $\beta X$111111 More precisely, such a retract descends to one of open sets $r^{-1}(U)\to U$; applying $\mathrm{H}^{q}(-;\mathcal{W})$ to the identity map $U\hookrightarrow r^{-1}(U)\to U$ then shows that if $n$ witnesses item 2 of Lemma 4.8 for $\beta X$ then it does for $T$ as well. Thus for deriving a contradiction, via Lemma 4.8, from the premises of Lemma 4.7, we may indeed restrict attention to the free $\mathsf{ED}$ spaces $\beta X$ for $X$ discrete. The characterization just alluded to is the correspondence, given by Stone duality, between open subsets $U$ of $\beta X$ and ideals $\mathcal{I}$ on $X$ [42, Lemma 19.1]. More precisely, any such $U$ equals $\bigcup_{I\in\mathcal{I}_{U}}N_{I}$ for some ideal $\mathcal{I}_{U}\subseteq P(X)$; note next that each such $N_{I}$ corresponds, via the map $\mathcal{U}\mapsto\mathcal{U}\cap P(I)$, to the set of ultrafilters on $I$, and that this map in fact defines a homeomorphism $N_{I}\to\beta I$. In consequence, for any field $K$ and vector space $W=\bigoplus_{\kappa}K$ and open $U\subseteq\beta X$, we have the following:
@@
Here the fourth and third equalities follow from the definition and the compactness of $\beta I$, respectively; the first is immediate from our discussion above. The second may be viewed as an instance of [68, Prop. 3.6.3]; more precisely (since directed colimits are exact functors on sheaf categories), the right hand side of line 13 is the $q^{\mathrm{th}}$ cohomology group of what may be written as
@@
where $\iota:\beta I\to U$ is the inclusion, the first equality is by way of the aforementioned reference, and the second is just as in the argument of Lemma 4.8. As noted above, the cohomology groups $\mathrm{Ext}^{r}(\mathcal{K}_{\beta I},\iota^{*}\mathcal{W}_{U})=\mathrm{H}^{r}(\beta I;\mathcal{W})$ of the last line’s interior term vanish in all degrees $r>0$. We arrive in this way to the expression in line 14 above.
The reward for these analyses is the expression appearing in line 16. Observe, for example, that if $\kappa=1$ in the sequence (13)–(16) then for any open $U\subseteq\beta X$ the $\mathcal{I}_{U}$-indexed inverse system of line 16 is plainly flasque, implying the overall expression zero for any $q>0$, from which it follows by the argument of Lemma 4.8 that the injective dimension of $\mathcal{K}$ over any $\mathsf{ED}$ space $S$ is at most one. Together with Lemma 4.14 below, this shows the first part of Theorem 4.3. The second part derives from Lemmas 4.7 and 4.8 (or, more precisely, from the argument of its (1$\Rightarrow$(2)) and the following, which we will argue via line 16 above as well.
@@
For any $n>0$, the associated notion of $n$-coherence is then that of a family
@@
$\mathrm{R}^{n}\mathrm{lim}_{I\in\tilde{\mathcal{I}}}\,\bigoplus_{\kappa}\prod_{I}K=0$
@@
By equations 13 through 16, it will suffice to show that for any $K$, $\kappa$, $X$, and $n$ as in the statement of the theorem, there exists an ideal $\mathcal{I}_{n}$ on $X$ such that $\mathrm{R}^{n}\mathrm{lim}_{I\in\mathcal{I}_{n}}\,\bigoplus_{\kappa}\prod_{I}K\neq 0$. For this it will suffice to exhibit for each $n>0$ a nontrivial $n$-coherent family of functions

### v2

@@ 4. Products of compact projectives
[…]
@@
    If $\mathcal{C}=\mathsf{Cond(Ab)}$ is the category of condensed abelian groups, then $\mathcal{C}^{\mathrm{cp}}$ is the category of direct summands of $\mathbb{Z}[S]$ for extremally disconnected $S$, which generates $\mathcal{C}$ under small colimits. [76, pp. 74-75]
@@
This proposition traces at the latest to Walter Rudin around 1960, by way of the attribution in [32], and induces, in turn, a number of effects and questions within the field of condensed mathematics. By item (4) above, for example, it implies that $\mathsf{Cond(Set)}^{\mathrm{cp}}$ is not closed under products. Subtler questions attach to item (5). There, $\mathbb{Z}[-]:\mathsf{Cond(Set)}\to\mathsf{Cond(Ab)}$ denotes the left adjoint to the forgetful functor $\mathsf{Cond(Ab)}\to\mathsf{Cond(Set)}$, and within $\mathsf{Cond(Ab)}$, the relation $\mathbb{Z}[S]\otimes\mathbb{Z}[T]=\mathbb{Z}[S\times T]$ renders $\mathbb{Z}[-]$-images of products of $\mathsf{ED}$ spaces computationally ubiquitous. These are not, in general, projective, although the main result in this direction requires more argument than one might at first expect; the proof given in [25], for example, answers an open question from [4] in Banach space theory [^10] That question, of whether the Banach space $C(\beta\mathbb{N}\times\beta\mathbb{N})$ is separably injective, typifies how this stratum of projectivity questions dualizes to injectivity questions in Banach settings; see [25, Prop. 3.17]. along the way:
@@
See [25, Appendix to III and pp. 26–27] for further discussion of these issues; the latter includes, for example, the following observation: by item (5), $\mathsf{Cond(Ab)}$ is rich in projective objects, i.e., in $A=\mathbb{Z}[S]$ for which $\mathrm{Ext}^{i}_{\mathsf{Cond(Ab)}}(A,-)=0$ for all $i>0$. For such to be internally projective, though, would entail that the internal Ext functors
@@
For any finite field $K$ and $\mathsf{ED}$ space $S$, the constant sheaf $\mathcal{K}$ on $S$ is injective. In contrast, if $\mathsf{ED}$ spaces $S$ and $T$ are each Čech-Stone compactifications of sets of cardinality at least $\aleph_{\omega}$, then for any field $K$ the injective dimension of the constant sheaf $\mathcal{K}$ on their product is infinite.
@@ 4.1. Sheaves on $\mathsf{ED}$ spaces and their products
We recall a few further fundamental points before turning to our main argument; for a fuller review, see, for example, [48, Chapter 2].
@@
Inferences about sheaves $\mathcal{G}$ on $X$ frequently reduce to inferences about their stalks $\mathcal{G}_{x}$ $(x\in X)$; relatedly, the sheafification of a presheaf $\mathcal{F}$ may be defined simply as that sheaf $\widetilde{\mathcal{F}}$ on $X$ which is equipped with a map $\mathcal{F}\to\widetilde{\mathcal{F}}$ inducing an isomorphism $\mathcal{F}_{x}\to\widetilde{\mathcal{F}}_{x}$ of each stalk.
@@
with the left-hand side computed, of course, in the category concluding the previous sentence [48, 2.7.5]. Therein, injective objects admit the standard definition; there are enough of them, and the injective dimension $\mathrm{injdim}_{X}(\mathcal{F})$ of any sheaf $\mathcal{F}$ is the minimal length of an injective resolution of $\mathcal{F}$.
A useful injectivity criterion is the following:
Lemma 4.7.
For any field $K$, a sheaf $\mathcal{F}$ of $K$-vector spaces over a topological space $X$ is injective if and only if $\mathrm{Ext}_{K}^{1}(\iota_{!}\mathcal{K},\mathcal{F})=0$ for every inclusion $\iota$ of a locally closed subset into $X$.
Proof.
The forward implication is standard. For the reverse implication, suppose that $\mathrm{Ext}_{K}^{1}$ vanishes in the manner described and let $f:\mathcal{A}\to\mathcal{F}$ and $g:\mathcal{A}\hookrightarrow\mathcal{B}$ be morphisms of sheaves over $X$; we need to show that $f$ extends to an $h:\mathcal{B}\to\mathcal{F}$. To that end, write $\mathcal{A}^{\prime}\leq\mathcal{A}^{\prime\prime}$ if $\mathcal{A}^{\prime}$ is a subsheaf of $\mathcal{A}^{\prime\prime}$, regard $\mathcal{A}$ as a subsheaf of $\mathcal{B}$, and consider the natural partial ordering of pairs $(\mathcal{A}^{\prime},h^{\prime})$ for which $\mathcal{A}\leq\mathcal{A}^{\prime}\leq\mathcal{B}$ and $h^{\prime}:\mathcal{A}^{\prime}\to\mathcal{F}$ satisfies $f=h^{\prime}\circ g$. Since this order satisfies the conditions of Zorn’s Lemma, our task reduces to showing that if there exists for some such $(\mathcal{A}^{\prime},h^{\prime})$ an $x\in S$ with $\mathcal{A}_{x}^{\prime}\lneq\mathcal{B}_{x}$ then there exists an $\mathcal{A}^{\prime\prime}$ with $\mathcal{A}^{\prime}\lneq\mathcal{A}^{\prime\prime}\leq\mathcal{B}$ and an $h^{\prime\prime}:\mathcal{A}^{\prime\prime}\to\mathcal{F}$ properly extending $h^{\prime}$.
[…]
$0\to\mathcal{A}^{\prime}\to\mathcal{A}^{\prime}+\varepsilon_{!}\mathcal{K}\to\iota_{!}\mathcal{K}\to 0$
of sheaves on $X$. Our premise, applied to the fragment of the induced long exact sequence
$\mathrm{Hom}_{K}(\mathcal{A}^{\prime}+\varepsilon_{!}\mathcal{K},\mathcal{F})\xrightarrow{e}\mathrm{Hom}_{K}(\mathcal{A}^{\prime},\mathcal{F})\to\mathrm{Ext}_{K}^{1}(\iota_{!}\mathcal{K},\mathcal{F}),$
implies that $e$ is surjective, and hence that $h^{\prime}$ extends to an an element $h^{\prime\prime}:\mathcal{A}^{\prime}+\varepsilon_{!}\mathcal{K}\to\mathcal{F}$ of $e^{-1}(h^{\prime})$. This $h^{\prime\prime}$ and $\mathcal{A}^{\prime\prime}=\mathcal{A}^{\prime}+\varepsilon_{!}\mathcal{K}$ conclude the argument. ∎
Let us see how this criterion implies the first part of Theorem 4.3 (compare also the proof of [38, Lemma VII.1.7]).
Lemma 4.8.
For any finite field $K$ and $\mathsf{ED}$ space $S$, the constant sheaf $\mathcal{K}$ on $S$ is injective.
Proof.
Observe first that since $K$ is compact, any continuous map from an open $U\subseteq S$ to $K$ uniquely extends to a continuous map $\beta U\to K$. Similarly, since $S$ is $\mathsf{ED}$, the inclusion $U\hookrightarrow S$ uniquely extends to an embedding of $\beta U$ as an open and closed subspace of $S$. Together these facts imply that the restriction map $\mathcal{K}(S)\to\mathcal{K}(U)$ is surjective. When the latter holds, as here, for all open $U\subseteq S$, we say that the sheaf in question is flasque (the term’s application to inverse systems above was a special instance); the interest of this property is that flasque resolutions are acyclic and therefore convenient for cohomology computations. It is easy to see that $\iota^{*}$ preserves this property for any such $\iota:U\hookrightarrow S$, and thus that $\mathrm{Ext}^{1}_{K}(\iota_{!}\mathcal{K},\mathcal{K})=\mathrm{H}^{1}(U;\iota^{*}\mathcal{K})=0$ for all such $\iota$ (the first equality follows from the $\iota_{!}$$\iota^{*}$ adjunction together with the fact that such $\iota^{*}$ are exact and preserve injectives [48, II.6]). By Lemma 4.7, it remains only to see that $\mathrm{Ext}^{1}_{K}(\iota_{!}\mathcal{K},\mathcal{K})=0$ holds for all inclusions $\iota$ of locally closed subsets $L$ into $S$ as well.
[…]
$0\to\eta_{!}\mathcal{K}\to\varepsilon_{!}\mathcal{K}\to\iota_{!}\mathcal{K}\to 0$
(cf. [48, p. 112]). Applying $\mathrm{Hom}_{K}(-,\mathcal{K})$ together with the reasoning above yields the long exact sequence fragment
$\mathrm{Hom}_{K}(\varepsilon_{!}\mathcal{K},\mathcal{K})\xrightarrow{f}\mathrm{Hom}_{K}(\eta_{!}\mathcal{K},\mathcal{K})\to\mathrm{Ext}_{K}^{1}(\iota_{!}\mathcal{K},\mathcal{K})\to 0.$
Flasqueness implies that $f:\mathcal{K}(U)\to\mathcal{K}(V)$ is surjective, and thus that
$\mathrm{Ext}^{1}_{K}(\iota_{!}\mathcal{K},\mathcal{K})=0.$
As $L$ was arbitrary, this completes the proof. ∎
For the second part of Theorem 4.3, we first prove a sequence of preliminary lemmas.
@@
of $\pi_{*}\mathcal{K}$ over $s$. By the compactness of $T$, for every locally constant $f:U\times T\to K$ there exists a neighborhood $V$ of $s$ and locally constant $g:T\to K$ such that $f(s^{\prime},t)=g(t)$ for all $(s^{\prime},t)\in V\times T$. It follows that $(\pi_{*}\mathcal{K})_{s}$ equals the $K$-vector space of continuous functions $T\to K$ (with discrete codomain); moreover, the latter is isomorphic to $W=\bigoplus_{w(T)}K$ by the argument of [39, Corollary 97.7]. The lemma’s concluding claim is now immediate: since the natural map from the constant sheaf $\mathcal{W}$ on $S$ to $\pi_{*}\mathcal{K}$ induces isomorphisms of stalks, it is an isomorphism of sheaves.
@@
for any $q\geq 0$, where $\iota^{*}$ denotes the pullback of the inclusion $\pi^{-1}(s)\hookrightarrow S\times T$ [66, Theorem 17.2]. Since pullback functors are exact and $\mathrm{H}^{q}(T;-)=0$ for any $q>0$ and profinite $T$ [87, Theorem 5.1], this in turn implies that $\pi_{*}$ is exact. The lemma’s first claim now follows from the general fact that pushforwards preserve injectives [48, II.4.1.3], since this, when combined with the foregoing, implies that the image $\pi_{*}\mathcal{J}$ of a finite injective resolution $\mathcal{J}$ of $\mathcal{K}$ is a finite injective resolution of $\pi_{*}\mathcal{K}$. ∎
@@
For every finite field $K$ and $\mathsf{ED}$ space $S$ and $K$-vector space $W$, the following conditions are equivalent:
@@
More exact relationships will emerge in the course of our proof below: the injective dimension of a $\mathcal{W}$ as in (1), for example, plays the role of an $n$ as in (2); an $n$ as in (2), on the other hand, implies that $\mathrm{injdim}_{S}(\mathcal{W})\leq n+1$. Note also that our assumptions on $K$ and $S$ are unneeded for the implication $(1)\Rightarrow(2)$.
@@
$(1)\Rightarrow(2)$: Much as in preceding arguments, the point is that the pullback functor $\iota^{*}$ associated to the inclusion $\iota:U\hookrightarrow S$ is both exact and preserves injectives. Hence $\iota^{*}$ applied to a length-$n$ injective resolution of the constant sheaf $\mathcal{W}$ over $S$ is a length-$n$ injective resolution of the constant sheaf $\mathcal{W}=\iota^{*}\mathcal{W}$ over $U$; this implies (2).
$(2)\Rightarrow(1)$: Fix an $n$ as in (2) and a resolution
(13) $0\to\mathcal{W}\to\mathcal{J}^{0}\to\cdots\to\mathcal{J}^{n-1}\to\mathcal{F}\to 0$
@@
of the sheaf $\mathcal{W}$ over $S$ in which each of the $\mathcal{J}^{i}$ are injective (if $n=0$, we just have the resolution $0\to\mathcal{W}\to\mathcal{F}\to 0$). By assumption, for every $q>n$ and inclusion $\iota$ of an open subset $U$ into $S$,
$0=\mathrm{H}^{q}(U;\mathcal{W})$ $=\mathrm{H}^{q}(U;\iota^{*}\mathcal{W})$
$=\mathrm{Ext}_{K}^{q}(\mathcal{K},\iota^{*}\mathcal{W})$
$=\mathrm{Ext}_{K}^{q}(\iota_{!}\mathcal{K},\mathcal{W})$
(14) $=\mathrm{Ext}_{K}^{q-n}(\iota_{!}\mathcal{K},\mathcal{F}).$
@@
Here the penultimate equality follows just as in Lemma 4.7 from the $\iota_{!}$$\iota^{*}$ adjunction together with the fact that such $\iota^{*}$ are exact and preserve injectives, and the last equality is via a standard dimension-shifting argument, such as appears, for example, in [86, Exercise 2.4.3]. The vanishing of line 4.1 for all $q\geq n$ and open $U\subseteq S$ doesn’t suffice for the conclusion of Lemma 4.7, but it’s enough in this context for something almost as good, namely that $\mathrm{injdim}_{S}(\mathcal{F})\leq 1$. To see this, fix a resolution
(15) $0\to\mathcal{F}\to\mathcal{J}^{n}\to\mathcal{H}\to 0$
with $\mathcal{J}^{n}$ injective. Application of $\mathrm{Hom}_{K}(\iota_{!}\mathcal{K},-)$ for open inclusions $\iota$ as above then induces long exact sequences telling us two things: (i) that $\mathrm{Ext}^{r}_{K}(\iota_{!}\mathcal{K},\mathcal{H})=0$ for all such $\iota$ and $r>0$, and (ii) that $\mathcal{H}$ inherits the flasqueness of $\mathcal{J}^{n}$. From these points, we conclude exactly as in Lemma 4.8 that $\mathcal{H}$ is injective. Splicing (15) to (13) along $\mathcal{F}$ then defines a length-$(n+1)$ injective resolution
$0\to\mathcal{W}\to\mathcal{J}^{0}\to\cdots\to\mathcal{J}^{n-1}\to\mathcal{J}^{n}\to\mathcal{H}\to 0$
of $\mathcal{W}$. ∎
@@
Against this background, we turn to the groups $\mathrm{H}^{q}(U;\mathcal{W})$ for such $W$ and open subsets $U\subseteq S$ as appear in Lemmas 4.9 and 4.10 above; since such $U$ admit a convenient characterization when $S$ is the Čech-Stone compactification $\beta X$ of some discrete space $X$, we will focus our attention on $S$ of this sort. Note that there is no loss of generality in doing so, since every $\mathsf{ED}$ space $T$ is a retract of some such $\beta X$.111111 More precisely, such a retract descends to one of open sets $r^{-1}(U)\to U$; applying $\mathrm{H}^{q}(-;\mathcal{W})$ to the identity map $U\hookrightarrow r^{-1}(U)\to U$ then shows that if $n$ witnesses item 2 of Lemma 4.10 for $\beta X$ then it does for $T$ as well. Thus for deriving a contradiction, via Lemma 4.10, from the premises of Lemma 4.9, we may indeed restrict attention to the free $\mathsf{ED}$ spaces $\beta X$ for $X$ discrete. The characterization just alluded to is the correspondence, given by Stone duality, between open subsets $U$ of $\beta X$ and ideals $\mathcal{I}$ on $X$ [45, Lemma 19.1]. More precisely, any such $U$ equals $\bigcup_{I\in\mathcal{I}_{U}}N_{I}$ for some ideal $\mathcal{I}_{U}\subseteq P(X)$; note next that each such $N_{I}$ corresponds, via the map $\mathcal{U}\mapsto\mathcal{U}\cap P(I)$, to the set of ultrafilters on $I$, and that this map in fact defines a homeomorphism $N_{I}\to\beta I$. In consequence, for any field $K$ and vector space $W=\bigoplus_{\kappa}K$ and open $U\subseteq\beta X$, we have the following:
@@
Here the fourth and third equalities follow from the definition and the compactness of $\beta I$, respectively; the first is immediate from our discussion above. The second may be viewed as an instance of [71, Prop. 3.6.3]; more precisely (since directed colimits are exact functors on sheaf categories), the right hand side of line 16 is the $q^{\mathrm{th}}$ cohomology group of what may be written as
@@
where $\iota:\beta I\to U$ is the inclusion, the first equality is by way of the aforementioned reference, and the second is just as in the arguments of Lemmas 4.8 and 4.10. As noted above, the cohomology groups $\mathrm{Ext}^{r}(\mathcal{K}_{\beta I},\iota^{*}\mathcal{W}_{U})=\mathrm{H}^{r}(\beta I;\mathcal{W})$ of the last line’s interior term vanish in all degrees $r>0$. We arrive in this way to the expression in line 17 above.
[…]
@@
Let $\mathbf{Y}[\kappa,X,\tilde{\mathcal{I}},K]$ denote the inverse system $\mathbf{X}[\mathcal{I},\mathcal{J},H]$ derived from this data as in Definition 2.16. Note that the set $\{\kappa\times I\mid I\in\tilde{I}\}$ is cofinal in $\mathcal{I}$ and that
$X_{\kappa\times I}=\bigoplus_{\kappa}\prod_{I}K$
for any $I\in\tilde{I}$. For any $n>0$, the associated notion of $n$-coherence is then that of a family
@@
$\operatorname{lim}^{n}\mathbf{Y}[\kappa,X,\tilde{I},K]=0$
@@
By equations 16 through 19, it will suffice to show that for any $K$, $\kappa$, $X$, and $n$ as in the statement of the theorem, there exists an ideal $\mathcal{I}_{n}$ on $X$ such that $\operatorname{lim}^{n}\mathbf{Y}[\kappa,X,\mathcal{I}_{n},K]\neq 0$. For this it will suffice to exhibit for each $n>0$ a nontrivial $n$-coherent family of functions